OpenStudy (anonymous):

Eliminate the parameter. x = 5t, y = t + 8 y = 5x + 8 y = x divided by five + 8 y = 5x - 8 y = x divided by five - 8

OpenStudy (anonymous):

x=5t, t=x/5 y=x/5 +8

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

how about in this one?

OpenStudy (anonymous):

Eliminate the parameter. x = square root of t, y = 3t + 7 y = 3x2 + 7 y = 3square root of x + 7, x ≥ 0 y = 3square root of x - 7, x ≥ 0 y = 3x2 - 7

OpenStudy (anonymous):

@surjithayer can u help me?

OpenStudy (anonymous):

\[x=\sqrt{t}\] \[x^{2}=t\] y=3t+7 \[y=3x ^{2}+7\]

OpenStudy (anonymous):

thank you! :)